OpenStudy (anonymous):

double check my answer: What is the value of the expression 5 – c – d when c = 8 and d = –13? A. –26 B. –16 C. 0 D. 10 my answer is a.

OpenStudy (anonymous):

The correct answer is B 5 -8 = -3. -3 - (-13) = -16 So B is your answer